OpenStudy (anonymous):

Use the definition of derivative to find the slope of the tangent of f at x=-1. Then find an equation of the tangent line. f(x) =3x^2

OpenStudy (raden):

m = f ' (1) can u find the derivative of f, first ?

OpenStudy (raden):

oppp.. m = f ' (-1), i mean

OpenStudy (anonymous):

So do i just plug -1 in as x in f(x)=3x^2?

OpenStudy (raden):

that's used too, the other componet is f ' (-1) so, find f ' first ?

OpenStudy (anonymous):

How do I find f ' ? this is in a math packet I have to do but we were never taught derivatives

OpenStudy (anonymous):

\[f'(-1)=\lim_{x\to -1}\frac{f(x)-f(-1)}{x+1}\]

OpenStudy (anonymous):

that is by the "definition"

OpenStudy (anonymous):

How do I use that to find the sloap of the tangent to f at x=-1? or it that it?

OpenStudy (anonymous):

\[\frac{3x^2-3}{x+1}=\frac{3(x+1)(x-1)}{x+1}=3(x-1)\] then replace \(x\) by \(-1\) and get \[3(-1-1)=-6\]

OpenStudy (anonymous):

slope*

OpenStudy (anonymous):

that is the slope, \(-6\)

OpenStudy (anonymous):

and the equation of the tangent line?

OpenStudy (anonymous):

you have the slope, it is \(-6\) and you have the point, it is \((-1,3)\) use the "point-slope" formula